find the product of the complex number and its conjugate. -3+2i
Answer by stanbon(75887)

find the product of the complex number and its conjugate. -3+2i
-----
Ans: (-3+2i)(-3-2i) = (-3)^2-(2i)^2 = 9+4 = 13
-----------------
Cheers,
Stan H.
----------------